# A solvable embedding mechanism for one-dimensional spinless and Majorana fermions in higher-dimensional spin-$\frac{1}{2}$ magnets

### v1

@@ Abstract
Stability to off-chain deformations in the first three cases of unfrustrated $XY$, frustrated $XY$ and $XX$ models is evidenced by the presence of the sector gap as shown in Figs. 24. This is because it is precisely these deformations such as off-chain $XY$ or $XX$ couplings that would lead to quantum fluctuations in the off-chain spin $Z$-values that were otherwise conserved by construction. Presence of the sector gap implies that if the energy scale associated with these fluctuations is lesser than the gap, then the ground states should be stable till they become large enough. Thus the magnetic states discussed in this paper are stable states of magnetic matter. In the case of the unfrustrated $XX$ model with an extensive degeneracy from the paramagnetic off-chain spins, the off-chain deformations would be a possible mechanism for the development of the conjectured slow mode.
Figure 7: Some imagined quantum architectures.
@@
Physical realization:– With a view towards physical realization, we can make the following remarks. In terms of materials, one may imagine an insulator material with localized $S=\frac{1}{2}$ moments with a crystal structure and quantum chemistry that accommodates a grid structure of easy-axis and easy-plane spin chains as the case may be. If the frustrated $XY$ case could be so realized, it would allow for highly anisotropic spin transport along the spin chains which can have practical applications such as in spin-based classical technologies. This could be an interesting avenue of exploration for the materials design and search community as alternative to the Kitaev material class with bond-dependent magnetic couplings.
@@
But perhaps the more natural application of the ideas laid out in this work may be in the realm of quantum technologies involving artificial platforms with controllable qubits. The grid structure proposed here can provide a blueprint for a bus architecture for quantum information traffic. The conserved off-chain spins can sequester the spin information to the on-chain spin chains. Some imagined architectures are shown in Fig. 7. For example the right panel of Fig. 7 could be an use case where turning on and off the dashed $ZZ$-couplings would act as a valve for spin information flow. Interestingly, such an architecture has been recently reported in Ref. [40] with a different application as its focus.
@@
Acknowledgements:– Funding support from SERB-DST, India via Grant No. MTR/2022/000386 and partially by Grant No. CRG/2021/003024 is acknowledged. I thank Sumilan Banerjee, Ganapathy Baskaran, Soumya Bera, Subhro Bhattacharjee, Hitesh Changlani, Kedar Damle, Nisheeta Desai, Giniyat Khaliullin, Gregoire Misguich, Laura Messio, Subroto Mukherjee, Ganpathy Murthy, Hridis Pal and Sumathi Rao for discussions. I thank Subhro in particular for the perturbative argument in the $J_{ZZ}/J_{XY}\ll 1$ limit. I am grateful to the Toulouse strongly correlated electron systems group, Laboratory of Theoretical Physics, University Paul Sabatier for the hospitality and discussions during a part of the work. I especially thank Fabien Alet for encouragement and feedback. I acknowledge the hospitality and support of ICTP, Trieste during a part of the work. I thank CEPIFRA for support during a part of the work.

### v2

@@ IV.1 Stability considerations
It is also important to consider the stability of the grounds states discussed here with respect to natural deformations of the Hamiltonian, i.e. off-chain XY couplings and on-chain Z-Ising couplings. For the XY variants with on-chain Z-Ising couplings as deformations, one can invoke Luttinger liquid arguments to argue for stability [38, 2]. The bipartite XY case would be stabler to on-chain Z-Ising deformations due to the presence of the gap in the on-chain spinless fermionic excitations. The same would be expected for the non-bipartite and bipartite X-Ising case due to the gap in the on-chain spectrum in presence of Ising order or disorder. Stability to off-chain deformations in the first three cases of bipartite XY, non-bipartite XY and X-Ising models is evidenced by the presence of the sector gap as shown in Figs. 35. This is because it is precisely these deformations such as off-chain XY or X-Ising couplings that would lead to quantum fluctuations in the off-chain spin Z-values that were otherwise conserved by construction. Presence of the sector gap suggests that if the energy scale associated with these fluctuations is lesser than the gap, then the ground states should be stable till they become large enough.
For the particular case of non-bipartite XY construction with gapless on-chain excitations, one may worry that the inter-chain interactions induced through the gapped off-chain excitations — by integrating out the off-chain spin degrees of freedom — may destabilize the system and lead to a gapped state for the on-chain spectrum as well [39, 40]. However, given the perturbative stability arguments in both $\frac{J_{\text{XY}}}{J_{\text{Z-I}}}\ll 1$ and $\frac{J_{\text{XY}}}{J_{\text{Z-I}}}\gg 1$ limits for the gapless on-chain state within the solvable bond-dependent parameter regime as discussed previously in Sec. III.4, we conjecture that additional off-chain perturbations will lead to a locking of the spin superfluid order parameter across the chains in the ground state. This suggests that in presence of off-chain perturbations a long-range ordered spin superfluid ground state – without a gap in the thermodynamic limit – obtains in the non-bipartite XY construction up to a finite range of such local perturbations, outside of which a gap may open. Even in the solvable regime, one may wonder if finite (low) temperatures can lead to a locking of the spin superfluid order across the chains via an order-by-disorder mechanism.
Note that the excitations on top of a long-ranged ordered spin superfluid ground state in two dimensions are bosonic magnons. Even so, the fermionic spinon excitation physics discussed in this work can play a role in a “dimensional reduction” sense, i.e. in an appropriate regime of temperature, the coupling ratio $\frac{J_{\text{XY}}}{J_{\text{Z-I}}}$ and other couplings (including an external magnetic field in Z direction) that take us out of the solvable regime, the effective description may still be in terms of on-chain spinon excitations and localized spin flip excitations with majority weight on the off-chain spins. A recent experimental example of a quasi-two-dimensional copper based quantum magnet showing such dimensional reduction phenomenology can be found in Ref. [41]. See also Ref. [42] in the context of two-dimensional $SU(2)$-symmetric Heisenberg magnets with magnonic excitations where “where a broad spectral-weight continuum at wave vector $q=(\pi,0)$ was interpreted as deconfined spinons” in relation to neutron-scattering experiments on another copper based compound [43].
Explicit verification of the above arguments may be done using various Quantum Monte Carlo or Matrix product based methods. In the case of the bipartite X-Ising model with an extensive degeneracy from the paramagnetic off-chain spins, the off-chain deformations would be a possible mechanism for the development of the conjectured slow mode and the associated restoration of the third law of thermodynamics.
IV.2 Physical realization
We end with some remarks with a view towards physical realization. Firstly, the models discussed here in the JW-mapped fermionic form are reminiscent of Falicov-Kimball models where also itinerant electrons or Bloch waves interact with static or localized electrons. They have been invoked to study a variety of phenomena such as mixed-valence and metal-insulator transitions. More details on these electronic phenomenan can be found in the following reviews [44, 45, 46, 47]. See this Ref. [48] for an example of Falicov-Kimball models with Majorana excitations. The models discussed in this paper realize aspects of Falicov-Kimball physics, however and it must be emphasized, in a non-electronic context and only after a non-local transformation. In other words, the spin-spin correlations encode Falicov-Kimball type effects involving the itinerant electrons in a highly non-linear way. Furthermore, the possible geometries allowed by the embedding mechanism laid down here admits on-site potentials that act as a “shared” background for sites on different chains that are somewhat unnatural in the electronic context that motivated the Falicov-Kimball model in the first place, i.e. “as a set of localized states centered at the sites of the metallic ions in the crystal” [49].
@@ IV.2 Physical realization
In a fully localized context without any itinerant electrons, one may imagine physical realizations such as an insulator material with localized $S=\frac{1}{2}$ moments with a crystal structure and quantum chemistry that accommodates a grid structure of easy-axis and easy-plane spin chains as the case may be. If the non-bipartite XY case could be so realized, it would allow for highly anisotropic spin transport along the spin chains which can have practical applications such as in spin-based classical technologies. This could be an interesting avenue of exploration for the materials design and search community as alternative to the Kitaev material class with bond-dependent magnetic couplings.
@@
But perhaps the more natural application of the ideas laid out in this work may be in the realm of quantum technologies involving artificial platforms with controllable qubits. The grid structure proposed here can provide a blueprint for a bus architecture for quantum information traffic. The conserved off-chain spins can sequester the spin information to the on-chain spin chains. Some imagined architectures are shown in Fig. 9. For example the right panel of Fig. 9 could be an use case where turning on and off the dashed Z-Ising couplings would act as a valve for spin information flow. Interestingly, such an architecture has been recently reported in Ref. [50] with a different application as its focus.
@@
Acknowledgements:– I thank Sumilan Banerjee, Ganapathy Baskaran, Soumya Bera, Subhro Bhattacharjee, Hitesh Changlani, Kedar Damle, Nisheeta Desai, Giniyat Khaliullin, Gregoire Misguich, Laura Messio, Subroto Mukherjee, Ganpathy Murthy, Hridis Pal and Sumathi Rao for discussions. I thank Subhro in particular for the perturbative argument in the $J_{\text{Z-I}}/J_{\text{XY}}\ll 1$ limit. I am grateful to the Toulouse strongly correlated electron systems group, Laboratory of Theoretical Physics, University Paul Sabatier for the hospitality and discussions during a part of the work. I especially thank Fabien Alet for encouragement and feedback. Funding support from SERB-DST, India via Grant No. MTR/2022/000386 and partially by Grant No. CRG/2021/003024 is acknowledged. I acknowledge the hospitality and support of ICTP, Trieste during a part of the work. I thank CEPIFRA during a part of this work for support to the “Indo-French workshop on Topology and Entanglement in Quantum Matter 2024” held at the University of Toulouse, France between June 3rd-7th, 2024.